SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

IFR Trickle Injection June 2003


Problems with last experience of Trickle Injection
(Old Forward Endcap)

  • Main problems were trips of LV PS (also some trips of the HV)

New Forward Endcap

  • New HV distribution and new LV PS!
  • During normal run we had a problem with LV PS.
  • Problems which could appear with Trickle Injection in June:

    • Trip of HV PS ---> mainly in external layers ---> reset from epics (as usual).
    • Trip of LV PS ---> mainly external layers could be the cause ---> reset from switches in EH. We will provide clear labels and instructions.
    • There is also a risk that we need to go in!

    Plan for Trickle Injection

    • IFR will start with external layers (13-16) OFF.
    • Monitoring of the behavior of HV (currents, trips...) for layers 1-12.
    • Monitoring of the behavior of LV PS.
    • We will procede step by step to turn ON external layers
      • ---> each time monitoring the behavior.
    • The idea is try to turn ON also external layers,
      • ---> but at this moment we don't know if it is possible to do.
    • We will have to monitor FEC occupancy versus time behavior, in order to check if dead FECs occour after the trickle injection.

Mirco Andreotti
Last modified: Thu Jun 19 14:41:19 PDT 2003